# Excitation Amplitude Sampling for Low Variance Electronic Structure on Quantum Computers

### v1

@@ II.1 Mixed energy functional
where |$\ket{\phi_{i}^{a}}$ represents the trial mean-field state with an electron removed from the orbital $i$ (occupied in the Hartree-Fock state) and excited to the unoccupied orbital $a$, with double and higher excitations represented similarly as |$\ket{\phi_{ij...}^{ab...}}$, and continuing until all electrons are promoted to particle states. This conservation of particle number via consideration of the excitation amplitudes from the Fermi vacuum truncates the full set of bit strings in the statevector to those with a Hamming weight given by the number of electrons.
@@ II.2 Configurational sampling via classical shadows
$M\geq\frac{\log(L)}{\epsilon^{2}}\max_{j}||\hat{O}_{j}||^{2}_{\text{shadow}},$ (11)
where the shadow norm ||.$||.||_{\text{shadow}}$ is dependent on the type of measurements taken as part of the protocol [26] and $L$ is the number of operators in the set $\{\hat{O}_{j}\}$ we aim to estimate, all with a precision better than $\epsilon$.
@@
This pseudo-channel can be analytically inverted for certain ensembles of unitaries, $\mathcal{U}$. In particular, in this work we select $\mathcal{U}$ to be the ensemble of operators over the $N$-qubit Clifford group, $\mathcal{C}_{N}$, whose unitary operators can be decomposed into primitive Clifford gates with a depth which grows only logarithmically with $N$ [96]. In particular, since the $N$-qubit Clifford group constitutes a 3-design of the full unitary group, this mapping can be analytically inverted by integrating over the full unitary group [26]. This enables a reconstruction of the original density operator as
@@
We can now define our operator set, $\{\hat{O}_{j}\}$, in order to extract the configurational amplitudes of interest, $c_{ij}^{ab}$ and $c_{0}$, from Eq. 9 in order to evaluate Eq. 10. These amplitudes are obtained as coherences of $\rho$, which we can extract akin to the approaches recently employed in QC-QMC approaches [71, 74]. To do this, rather than working with the density operator $\rho$, we construct $\tilde{\rho}=|\tau\rangle\langle\tau|$, defined by
@@
Given the samples of the random Clifford operators, $U_{k}$, the matrix elements $\langle\phi_{ij}^{ab}|U_{k}^{\dagger}|b_{k}\rangle$ can be efficiently evaluated on classical computers by using the stabilizer formalism, requiring a classical time complexity of $\mathcal{O}[N^{3}]$; for further details see App. C [97]. Furthermore, due to the non-Gaussian nature of the sampled distribution in Eq. 12, it has been shown that a ‘median-of-means’ approach leads to a more effective bound on the expectation value than a simple mean [26]. Therefore in this work we always consider the expectation value as a median, after dividing the samples into three groups and finding the mean of each group.
The classical shadow formulation therefore allows for the joint measurement of all $\mathcal{O}[N^{4}]$ operators given in Eq. 15, from which the $c_{ij}^{ab}$ amplitudes required in Eq. 10 can be computed, by using the same set of $M$ samples of $\{U_{k},|b_{k}\rangle\}$. The precision of these resulting estimates depends on the ‘shadow norm’ of the operators under the ensemble $\mathcal{U}=\mathcal{C}_{N}$ in Eq. 11, which can be considered as the maximum variance of the expectation values over the set of all possible states, denoted by $||\hat{O}_{j}||^{2}_{\textrm{shadow}}$. For the $\mathcal{C}_{N}$ ensemble, the shadow norm for the operators of interest in Eq. 15 is bounded by their Hilbert-Schmidt norm, $||\hat{O}_{j}||^{2}_{\textrm{shadow}}\leq 3\mathrm{Tr}[\hat{O}_{j}^{2}]=6$. This norm is independent of $N$, and therefore we expect the precision by which we can extract any single $c_{ij}^{ab}$ amplitude to be independent of the system size. From Eq. 11, this results in the scaling in the number of shots for a maximum additive error in any amplitude over the entire set of required $c_{ij}^{ab}$ amplitudes ($\epsilon_{\text{amp}}$) as
@@ II.4 Numerical performance of mixed estimator
Figure 1: Variance of mixed and Pauli energy estimators for Hydrogen chains over changing correlation strengths and system size. (a) Mean and standard deviation of the energy estimate distributions of the Pauli and mixed estimators sampled with 1000 shots, for a ten atom symmetrically stretched Hydrogen system in an STO-3G basis. Also included is the exact solution (FCI, red) and coupled-cluster with singles and doubles (CCSD, navy) results, which failed to converge for distances above $1.5$Å. Both the Pauli and mixed estimators are obtained from samples of the exact ground statevector, with the Pauli estimate fully grouping the Pauli operators into commuting sets using a graph coloring algorithm, and the mixed estimate employing the Clifford shadow protocol as described in Sec. II.2. (b) Standard deviations of the Pauli (dashed) and mixed (solid) energy estimator distributions as the Hydrogen chain system size increases for a fixed interatomic distance of $1.5$Å, and a fixed 1000 shots per sample of the distribution. The black line again shows the Pauli operators fully grouped into as few commuting sets as possible, while the purple results group Pauli operators based on qubitwise commutivity. The orange line represents the standard deviations of the mixed estimator with the Hartree–Fock trial state using the Clifford shadows method, and is shown to exhibit approximately linear scaling of the error with system size.
@@
We compare the energy estimators applied to the symmetric stretching coordinate of a linear chain of Hydrogen atoms, without breaking any spatial or spin symmetries (i.e. a restricted basis). This is an important and widespread benchmark model for algorithmic development in ab initio systems. It features the full complexity of a realistic Hamiltonian with a correlation strength which can be tuned from weakly correlated at compressed geometries to strongly correlated at stretched cases [104, 37]. Figure 1(a) shows the potential energy over the stretching of the chain for $10$ atoms (20 qubits) in a minimal basis, comparing the energy and its standard deviation for the mixed and Pauli estimators while the number of shots is fixed at $1000$ for both methods. In addition, two reference classical methods are shown, with full configuration interaction (FCI) giving the exact solution, and coupled-cluster with single and double excitations (CCSD), providing an inexact yet classically tractable ansatz for comparison. Since CCSD is a resummed perturbative expansion, it fails for strongly correlated systems. Therefore, it provides excellent agreement with FCI for compressed geometries, but fails to converge on stretching beyond a bond length of $1.5$Å, which can be considered as the point at which strong correlations lead to the breakdown of standard perturbative approaches.
@@
The Pauli and mixed energy estimates are sampled from the exact ground state via statevector simulation. The Pauli estimator is given as the mean and standard deviation of the distribution of 1000 shot estimates (with the estimated standard deviation taken as the average of 100 simulations to lower the uncertainty in the true distribution). Commuting Pauli strings from the Jordan-Wigner encoding of the Hamiltonian are fully grouped using a heuristic graph coloring algorithm to find the optimal cliques [20]. This represents the standard approach to sampling the energy, with the version of this procedure restricted to qubitwise commuting groups the default in Qiskit Aer 0.16.1. The mixed estimator results are found with the same number of shots, but via the classical shadow procedure of Sec. II.2. We note that it is also possible to directly sample the expectation values of the operators of Eq. 15 required for the mixed estimator thereby avoiding the requirement of the classical shadow protocol entirely. Results from this are shown in App. D, and demonstrate the advantage which comes from the combination of both the classical shadow protocol with the mixed estimator.
While both sampled energy expectation values in Fig. 1(a) are numerically unbiased compared to the FCI reference, the results show that the Pauli estimator has an approximately constant standard deviation over the different bond lengths. This reflects that the variance in Eq. 4 is not explicitly dependent on the strength of the correlations in the system, although there can be indirect effects through the variance of the individual Pauli operators. This is in contrast to the mixed estimator, where the estimate in Eq. 5 has an explicit dependence on the overlap with the trial state |$\ket{\phi}$, which here is the Hartree–Fock state. As a result, as the chain is stretched, the variance of the energy estimate increases, as strong correlations reduce the overlap with the trial state to small values. Nevertheless, it is only at the final point in the near-dissociated atomic limit where we find the mixed estimator to exhibit a larger variance in the energy estimate than the Pauli estimator. Around equilibrium geometries of $1.0$Å, the standard deviation of the energy distribution with 1000 shots is nearly an order of magnitude smaller. Due to the $\epsilon^{-2}$ scaling in Eqs. 4 and 19, this leads to a more than 50-fold reduction in the number of shots required to extract the energy with a desired error when using the mixed compared to the Pauli estimator here. The reduction in the statistical error for the mixed estimator persists well into the regime where the chain is stretched beyond equilibrium, including the non-perturbative regime beyond $1.5$Å  where CCSD fails to converge. The estimators exhibit similar errors only in highly correlated settings once the overlap with the trial state has dropped below $0.25$.
@@
We can also consider the scaling of the error with system size. The naive arguments of Sec. II.3 suggest a scaling in the number of shots as a function of system size for a constant error is $\mathcal{O}[N\log(N)]$ for the mixed estimator, while up to potentially $\mathcal{O}[N^{5}]$ for the Pauli estimator (with the extra factor of $N$ arising from the empirical scaling of the Hamiltonian-dependent prefactors shown in App. B). However, this ignores the implicit exponential scaling of the mixed estimator due to the decreasing overlap with the trial state as the system grows, as discussed above. Figure 1(b) therefore shows the standard deviation of the two energy estimates for samples of $1000$ shots while increasing the number of atoms in the chain. The interatomic spacing is fixed at $1.5$Å  for all chain lengths – the point at which CCSD starts to fail for the ten atom system, and with the weight of the Hartree–Fock configuration $\approx 0.6$ for 12 atoms, indicating a moderately strongly correlated state.
@@ II.5 Noisy and approximate states
Figure 2: The effect of noise on the energy estimators for a fixed, optimized VQE ansatz, showing improved resilience of the mixed estimator. Energies for a four atom Hydrogen chain, computed via the Pauli (purple) and mixed (orange) estimators for different noise levels, each using eight sets of 1000 shots. The green points represent execution on the IBM Brisbane quantum device with the same number of shots, largely validating the performance of the simulated results at $\alpha=1$. The exact energy (FCI, dashed red) and Hartree–Fock energy (dashed light blue) are shown for comparison. The noise model for a given $\alpha$ is shown in Eq. 20, with $\alpha=0$ a noiseless simulation, and $\alpha=1$ approximating the noise of the real quantum device.
To gauge the applicability to realistic quantum devices, we consider the resilience of the mixed energy estimate with respect to noise and sampling of inexact states. While Eq. 5 must agree with the exact energy when the state |$\ket{\Psi}$ is an eigenstate of the Hamiltonian, it will not in general agree with the Pauli estimate of Eq. 3 for other states. We therefore consider the ground state of a four atom (8 qubit) Hydrogen chain obtained from the Variational Quantum Eigensolver (VQE), using the Local Unitary Cluster Jastrow (LUCJ) ansatz [105] with $23$ parameters, implemented and optimized using the ffsim package [106]. The restriction of the state to the employed ansatz leads to an inexact state with an energy error of $\sim 5\textrm{mE}_{h}$ after optimization. The optimized circuit was compiled for the IBMQ architecture and simulated with a custom noise model constructed to match the publicly available error rates for the IBM Eagle device. The model consisted of the $T_{1}$ and $T_{2}$ times, readout error, and one and two qubit gate error, assumed to be purely in the depolarising channel, giving an overall noise channel, $\mathcal{M}$. This error model was then linearly interpolated via a single parameter, $\alpha$, to the identity channel ($\mathcal{I}$) as
@@
In Fig. 2 we plot the energy of this system estimated from eight sets of 1000 shots, as the noise model is varied between $0\leq\alpha\leq 1$, to determine the mean and the error bar over the groups. We plot the Hartree-Fock energy (light blue) and exact FCI energy (red) for comparison and the results from the mixed estimator (orange) and the default IBM Pauli estimator for device execution (purple). We also validate the results for the full noise model ($\alpha=1$) by using the physical IBM Brisbane quantum device (green), which has the same processor architecture as the Eagle device used for the noise model. All measurements are taken from the same fixed VQE circuit (optimized via statevector simulation in the absence of noise) defining the state of interest. The physical quantum hardware results qualitatively agree with our noisy statevector simulations, validating the noise model employed.
@@
We find that as $\alpha$ increases from the noiseless simulation, the error in the Pauli estimator increases rapidly, such that the standard approach to estimating the energy as $\mathrm{Tr(\rho\mathcal{H})}$ is worse than the Hartree–Fock energy, even with $\alpha$ as low as $\sim 0.05$. As the noise continues to grow, the estimate becomes significantly worse than even a mean-field description of the electronic structure. This is due to an accumulation of errors in the circuit, despite the deliberately shallow LUCJ state ansatz, with all errors necessarily raising the energy variationally as the error rate increases. In contrast, the mixed estimator maintains an average much closer to the true value of the noiseless state, with increasing noise levels reflected primarily in a larger variance in the estimate, with a degree of error cancellation afforded by the non-variationality of the approach. In the large noise limit, the state will have a random overlap with the trial state in Eq. 8, providing an estimate given by the trial state energy (Hartree–Fock) plus some random change whose variance grows as the overlap decreases due to the noise. However, this effectively bounds the systematic error by the Hartree–Fock energy, and it is clear that there is a range of noise with $\alpha\lesssim 0.5$ where the mixed energy result remains a good approximation to the FCI result, and provides a statistically significant portion of the correlation energy commensurate with the LUCJ description. This is despite the formal Pauli estimator providing an energy substantially worse than the mean-field result, and therefore negative correlation energies.
@@
It has been noted in other contexts that estimates of these ratios of overlaps required for the mixed estimator benefit from an inherent resilience to noise, both in the shadow sampling procedure [73], and to global depolarizing noise [71, 74], which can often be considered a reasonable approximation to the cumulative effect of local noise [107, 108]. This inherent resilience in the mixed estimator means that a reduction in the noise of the quantum device by as little as $2-3$ times could allow a reasonable energy estimate to be obtained without error mitigation in many cases. This provides the possibility of extracting faithful correlation energies in noise regimes an order of magnitude larger than those required for meaningful Pauli energy estimates. While this partially obviates the need for error mitigation methods, their combination within the mixed estimator to further improve noise resilience remains an interesting route for further development [83]. However, as noted before, a drawback of the mixed energy functional is that it is unbounded, and therefore cannot be used to directly optimize a variational ansatz e.g. within VQE. Nevertheless, the significantly enhanced robustness to noise will still likely make it an attractive proposition in estimating the energy of already obtained states, both in the final step of a VQE workflow, and within other quantum algorithms.

### v2

@@ II.1 Mixed energy functional
[…]
@@ II.2 Configurational sampling via classical shadows
[…]
where the shadow norm $||.||_{\text{shadow}}$ is dependent on the type of measurements taken as part of the protocol [26] and $L$ is the number of operators in the set $\{\hat{O}_{j}\}$ we aim to estimate, all with a precision better than $\epsilon$.
@@
This pseudo-channel can be analytically inverted for certain ensembles of unitaries, $\mathcal{U}$. In particular, in this work we select $\mathcal{U}$ to be the ensemble of operators over the $N$-qubit Clifford group, $\mathcal{C}_{N}$, whose unitary operators can be decomposed into primitive Clifford gates with a depth which grows only logarithmically with $N$, even without all-to-all connectivity [96, 97]. This is likely far shallower than any electronic structure state preparation circuits, while further optimization and approximation of these Clifford unitaries to low gate counts is an active research area [98]. Importantly for their application to classical shadows, since the $N$-qubit Clifford group constitutes a 3-design of the full unitary group [99], this mapping can be analytically inverted by integrating over the full unitary group [26]. This enables a reconstruction of the original density operator as
@@
The benefits of classical shadows have spurred its combination with a number of approaches in the field of electronic structure, sampling from the local Clifford [28, 78], matchgate [30, 72, 100] or global Clifford operator sets [71, 101, 74] to jointly evaluate expectation values. In comparison to other approaches using global Clifford shadows, in this work we aim to efficiently evaluate the energy (and other expectation values) of an eigenstate already prepared on a quantum device, in contrast to the use of the shadow state from a quantum device to facilitate a classical simulation, for example in defining trial states for suppressing the sign problem in (classically computed) Monte Carlo methods [71, 74] or to evaluate matrix elements for a classically tractable subspace expansion of the low-energy physics [101].
We can now define our operator set, $\{\hat{O}_{j}\}$, in order to extract the configurational amplitudes of interest, $c_{ij}^{ab}$ and $c_{0}$, from Eq. 9 in order to evaluate Eq. 10. These amplitudes are obtained as coherences of $\rho$, which we can extract akin to the approaches recently employed in QC-QMC approaches mentioned previously [71, 74]. To do this, rather than working with the density operator $\rho$, we construct $\tilde{\rho}=|\tau\rangle\langle\tau|$, defined by
@@
Given the samples of the random Clifford operators, $U_{k}$, the matrix elements $\langle\phi_{ij}^{ab}|U_{k}^{\dagger}|b_{k}\rangle$ can be efficiently evaluated on classical computers by using the stabilizer formalism, requiring a classical time complexity of $\mathcal{O}[N^{3}]$; for further details see App. D [102]. Furthermore, due to the non-Gaussian nature of the sampled distribution in Eq. 12, it has been shown that a ‘median-of-means’ approach leads to a more effective bound on the expectation value than a simple mean [26]. Therefore in this work we always consider the expectation value as a median, after dividing the samples into three groups and finding the mean of each group.
The classical shadow formulation therefore allows for the joint measurement of all $\mathcal{O}[N^{4}]$ operators given in Eq. 15, from which the $c_{ij}^{ab}$ amplitudes required in Eq. 10 can be computed, by using the same set of $M$ samples of $\{U_{k},|b_{k}\rangle\}$. The precision of these resulting estimates depends on the ‘shadow norm’ of the operators under the ensemble $\mathcal{U}=\mathcal{C}_{N}$ in Eq. 11, which can be considered as the maximum variance of the expectation values over the set of all possible states, denoted by $\hat{O}_{j}||^{2}_{\textrm{shadow}}$. For the $\mathcal{C}_{N}$ ensemble, the shadow norm for the operators of interest in Eq. 15 is bounded by their Hilbert-Schmidt norm, $\hat{O}_{j}||^{2}_{\textrm{shadow}}\leq 3\mathrm{Tr}[\hat{O}_{j}^{2}]=6$. This norm is independent of $N$, and therefore we expect the precision by which we can extract any single $c_{ij}^{ab}$ amplitude to be independent of the system size. From Eq. 11, this results in the scaling in the number of shots for a maximum additive error in any amplitude over the entire set of required $c_{ij}^{ab}$ amplitudes ($\epsilon_{\text{amp}}$) as
@@ II.4 Numerical performance of mixed estimator
Figure 1: Variance of mixed and Pauli energy estimators for Hydrogen chains over changing correlation strengths and system size. (a) Mean and standard deviation of the energy estimate distributions of the Pauli and mixed estimators sampled with 1000 shots, for a ten atom symmetrically stretched Hydrogen system in an STO-3G basis. Also included is the exact solution (FCI, red) and coupled-cluster with singles and doubles (CCSD, navy) results, which failed to converge for distances above 1.51.5Å. Both the Pauli and mixed estimators are obtained from samples of the exact ground statevector, with the Pauli estimate fully grouping the Pauli operators into commuting sets using a graph coloring algorithm, and the mixed estimate employing the Clifford shadow protocol as described in Sec. II.2. (b) Standard deviations of the Pauli (dashed) and mixed (solid) energy estimator distributions as the Hydrogen chain system size increases for a fixed interatomic distance of 1.51.5Å, and a fixed 1000 shots per sample of the distribution. The black line again shows the Pauli operators fully grouped into as few commuting sets as possible, while the purple results group Pauli operators based on qubitwise commutivity. The orange line represents the standard deviations of the mixed estimator with the Hartree–Fock trial state using the Clifford shadows method, and is shown to exhibit approximately linear scaling of the error with system size.
@@
We compare the energy estimators applied to the symmetric stretching coordinate of a linear chain of Hydrogen atoms, without breaking any spatial or spin symmetries (i.e. a restricted basis). This is an important and widespread benchmark model for algorithmic development in ab initio systems. It features the full complexity of a realistic Hamiltonian with a correlation strength which can be tuned from weakly correlated at compressed geometries to strongly correlated at stretched cases [109, 37]. Figure 1(a) shows the potential energy over the stretching of the chain for $10$ atoms (20 qubits) in a minimal basis, comparing the energy and its standard deviation for the mixed and Pauli estimators while the number of shots is fixed at $1000$ for both methods. In addition, two reference classical methods are shown, with full configuration interaction (FCI) giving the exact solution, and coupled-cluster with single and double excitations (CCSD), providing an inexact yet classically tractable ansatz for comparison. Since CCSD is a resummed perturbative expansion, it fails for strongly correlated systems. Therefore, it provides excellent agreement with FCI for compressed geometries, but fails to converge on stretching beyond a bond length of 1.51.5Å, which can be considered as the point at which strong correlations lead to the breakdown of standard perturbative approaches.
@@
The Pauli and mixed energy estimates are sampled from the exact ground state via statevector simulation. This ensures that there is no bias in either the mixed estimator or Pauli estimator results, and therefore we can compare the variance of the estimators in isolation. The Pauli estimator is given as the mean and standard deviation of the distribution of 1000 shot estimates (with the estimated standard deviation taken as the average of 100 simulations to lower the uncertainty in the true distribution). Commuting Pauli strings from the Jordan-Wigner encoding of the Hamiltonian are fully grouped using a heuristic graph coloring algorithm to find the optimal cliques [20]. This represents the standard approach to sampling the energy, with the version of this procedure restricted to qubitwise commuting groups the default in Qiskit Aer 0.16.1. The mixed estimator results are found with the same number of shots, but via the classical shadow procedure of Sec. II.2. We note that it is also possible to directly sample the expectation values of the operators of Eq. 15 required for the mixed estimator thereby avoiding the requirement of the classical shadow protocol entirely. Results from this are shown in App. E, and demonstrate the advantage which comes from the combination of both the classical shadow protocol with the mixed estimator.
[…]
@@
We can also consider the scaling of the error with system size. The naive arguments of Sec. II.3 suggest a scaling in the number of shots as a function of system size for a constant error is $\mathcal{O}[N\log(N)]$ for the mixed estimator, while up to potentially $\mathcal{O}[N^{5}]$ for the Pauli estimator (with the extra factor of $N$ arising from the empirical scaling of the Hamiltonian-dependent prefactors shown in App. B). However, this ignores the implicit exponential scaling of the mixed estimator due to the decreasing overlap with the trial state as the system grows, as discussed above. Figure 1(b) therefore shows the standard deviation of the two energy estimates for samples of $1000$ shots while increasing the number of atoms in the chain. The interatomic spacing is fixed at 1.51.5Å  for all chain lengths – the point at which CCSD starts to fail for the ten atom system, and with the weight of the Hartree–Fock configuration $\approx 0.6$ for 12 atoms, indicating a moderately strongly correlated state.
@@ II.5 Noisy and approximate states
Figure 2: The effect of noise on the energy estimators for a fixed, optimized VQE ansatz, showing improved resilience of the mixed estimator. (a) Energies for a four atom Hydrogen chain, computed via the Pauli (purple) and mixed (orange) estimators for different noise levels, each using eight sets of 1000 shots. The points are the mean energy over these sets, characterizing the bias in the estimators due to the noise and inexact state, while the error bars characterize the standard error over these sets for the different estimators, indicating the statistical variance of the results due to finite shot numbers. The green points represent execution on the IBM Brisbane quantum device with the same number of shots, somewhat validating the performance of the simulated results at $\alpha=1$. The exact energy (FCI, dashed red) and Hartree–Fock energy (dashed light blue) are shown for comparison. The noise model for a given $\alpha$ is shown in Eq. 20, with $\alpha=0$ a noiseless simulation, and $\alpha=1$ approximating the noise of the real quantum device. (b) The mean-squared error (MSE) for the different estimators compared to the exact energy, incorporating both the bias and variance of the estimators in a single metric.
[…]
The optimized circuit was compiled for the IBMQ architecture with a custom noise model constructed to best reproduce the publicly available physical device error metrics. These included using the $T_{1}$ and $T_{2}$ relaxation times to parameterize a single-qubit thermal relaxation channel introducing phase and amplitude damping, one- and two-qubit gate error rates used to inject local depolarization, and local read out error rates designed to match the the IBM Eagle device, giving an overall noise channel, $\mathcal{M}$. This error model was then linearly interpolated via a single parameter, $\alpha$, to the identity channel ($\mathcal{I}$) as
@@
In Fig. 2 we plot the energy of this system estimated from eight sets of 1000 shots, as the noise model is varied between $0\leq\alpha\leq 1$, to determine the mean and the error bar over the groups. We plot the Hartree-Fock energy (light blue) and exact FCI energy (red) for comparison and the results from the mixed estimator (orange) and the default IBM Pauli estimator for device execution (purple). We also validate the results for the full noise model ($\alpha=1$) by using the physical IBM Brisbane quantum device (green), which has the same processor architecture as the Eagle device used for the noise model. All measurements are taken from the same fixed VQE circuit (optimized via statevector simulation in the absence of noise) defining the state of interest. The physical quantum hardware results qualitatively agree with our noisy statevector simulations, validating the noise model employed. In the right hand plot, we also show the mean squared energy error (MSE) compared to the exact results, combining the effects of both the bias and the variance in the estimators for this inexact energy, and demonstrating the combined benefit of the mixed estimator.
@@
We find that as $\alpha$ increases from the noiseless simulation, the bias in the Pauli estimator increases rapidly, such that the standard approach to estimating the energy as $\mathrm{Tr(\rho\mathcal{H})}$ is worse than the Hartree–Fock energy, even with $\alpha$ as low as $\sim 0.05$. As the noise continues to grow, the estimate becomes significantly worse than even a mean-field description of the electronic structure. This is due to an accumulation of errors in the circuit, despite the deliberately shallow LUCJ state ansatz, with all errors necessarily raising the energy variationally as the error rate increases. In contrast, the mixed estimator maintains a less biased estimate, with the average much closer to the true value of the noiseless state, with increasing noise levels reflected primarily in a larger variance in the estimate, with a degree of error cancellation afforded by the non-variationality of the approach. In the large noise limit, the state will have a random overlap with the trial state in Eq. 8, providing an estimate given by the trial state energy (Hartree–Fock) plus some random change whose variance grows as the overlap decreases due to the noise. However, this effectively bounds the systematic error by the Hartree–Fock energy, and it is clear that there is a range of noise with $\alpha\lesssim 0.5$ where the mixed energy result remains a good approximation to the FCI result, and provides a statistically significant portion of the correlation energy commensurate with the LUCJ description. This is despite the formal Pauli estimator providing an energy substantially worse than the mean-field result, and therefore negative correlation energies.
@@
It has been noted in other contexts that estimates of these ratios of overlaps required for the mixed estimator benefit from an inherent resilience to noise, both in the stochastic shadow sampling procedure [73], and to global depolarizing noise [71, 74], which can often be considered a reasonable approximation to the cumulative effect of local noise [112, 113]. This is because global depolarizing noise will exactly cancel in the ratio of probability amplitudes [71], and therefore leave the mixed estimator unchanged. While our simulated noise model goes well beyond just global depolarization, this inherent resilience in the mixed estimator means that a reduction in the noise of the quantum device by as little as 232-3 times could allow a reasonable energy estimate to be obtained without error mitigation in many cases. This provides the possibility of extracting faithful correlation energies in noise regimes an order of magnitude larger than those required for meaningful Pauli energy estimates. While this partially obviates the need for error mitigation methods, their combination within the mixed estimator to further improve noise resilience remains an interesting route for further development [83]. However, as noted before, a drawback of the mixed energy functional is that it is unbounded, and therefore cannot be used to directly optimize a variational ansatz e.g. within VQE. Nevertheless, the significantly enhanced robustness to noise will still likely make it an attractive proposition in estimating the energy of already obtained states, both in the final step of a VQE workflow, and within other quantum algorithms.